# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ XIV Global Sensing: Parameter Estimation with Minimal Prior Information
In contrast to local estimation theory, global estimation theory applies when no (or some) prior information is available. Recently, the impact of uncertainty in control parameters on the sensitivity of critical sensors has been explored Mihailescu et al. (2024c). Therefore, in these situations the unknown parameter varies within a broader sensing interval. Consequently, the choice of measurement to be performed on the probe and the analysis of the estimation data should be optimal on average, that is it should be equally effective for any possible value of the parameter Helstrom (1976); Holevo (2011a).
One of the earliest studies on global sensing, aimed at overcoming the limitations of local estimation theory, was proposed by van Trees van Trees (1968); Gill and Levit (1995) (a stronger type of inequality is also provided by Klaassen Klaassen (1989)). The van Trees inequality addresses the comprehensive concept of global sensing, where the unknown parameter $\lambda$ varies according to some prior distribution $z(\lambda)$, determining that the average variance is given by the Bayesian Cramér-Rao bound or posterior Cramér-Rao bound van Trees (1968); Paris (2009):
@@
Note that, in general, van Trees bound is not tight and cannot be saturated. To achieve a tighter bound a systematic and general metric evaluating the performance of global sensing and its significance for quantum many-body probes with criticality has been examined in Ref. Montenegro et al. (2021). The global sensing procedure involves considering a modified quantum Cramér-Rao bound, which introduces the average uncertainty of the estimation as follows:
@@
where $z(\lambda)$ is the prior distribution of the unknown parameter $\lambda$ to be sensed, $M$ is the number of measurements, $\mathcal{I}_{Q}(\lambda|\bm{B})$ is the QFI, $\bm{B}{=}(B_{1},B_{2},\ldots)$ are external tunable parameters interacting with the probe, and the unknown parameter varies over a sensing interval
@@
$H=J\sum_{i=1}^{L}\sigma_{x}^{i}\sigma_{x}^{i+1}-\sum_{i=1}^{L}(B_{z}+\lambda_{z})\sigma_{z}^{i},$ (90)
@@
where $L$ is the system size, $\sigma_{\alpha}^{i}$ ($\alpha{=}x,z$) represents the Pauli operator at site $i$, $J>0$ denotes the exchange interaction, $B_{z}$ corresponds to the controllable magnetic field that can be adjusted, $\lambda_{z}$ stands for the field to be estimated, and periodic boundary conditions are imposed. Assuming that the unknown parameter varies over a wide region $\Delta\lambda_{z}$, the protocol introduced in Ref. Montenegro et al. (2021) systematically optimizes the quantum many-body probe to deliver the best sensitivity performance on average. In the case of single-parameter estimation, the control field $B_{z}$ acts as an offset of the critical point, see Fig. 19(a). However, when the phase diagram becomes more complex or in the presence of multi-parameter estimation, such an offset becomes non-trivial, see Fig. 19(b).
@@
Figure 19: Sketch of phase diagram and the role of global sensing: $(a)$ The global sensing scheme tunes the quantum probe to operate optimally on average. This procedure ensures that the control fields will optimize the probe to exploit the critical point. $(b)$ As the phase diagram becomes more complex, the probe’s offset becomes non-trivial. In general, the minimization of $g(\bm{B})$ (see Eq. (87)) systematically optimizes the probe to exploit criticality across the entire phase diagram. Figure taken from Montenegro et al. (2021).
@@
By defining the size of the probe $L$ as the relevant sensing resource, it has been demonstrated that the QFI scales super-linearly with $L$ over a moderate wide interval of $\Delta\lambda_{z}$ Montenegro et al. (2021). Fig. 20(a) shows the fitting coefficients $a$ and $b$ obtained from the fitting function $g(B_{z}^{*},\Delta\lambda_{z})\sim aL^{-b}+c$, as a function of the interval $\Delta\lambda_{z}/J$. Here, $\Delta\lambda_{z}/J{\rightarrow}0$ represents local sensing. As the figure illustrates, the $b$ coefficient decreases from the (expected) Heisenberg limit ($b=2$) to the standard limit ($b=1$) as the interval $\Delta\lambda_{z}$ increases. The global sensing protocol still demonstrates quantum-enhanced sensitivity for sensing intervals $0\leq\Delta\lambda_{z}\lesssim 0.07J$. In Fig. 20(b), even under the seemingly standard limit (i.e., $b=1$), the global sensing strategy still delivers superior sensing precision compared to a non-optimized probe. This superiority arises from the global sensing strategy, which ensures probe optimization by minimizing $g(\bm{B})$ over the prior distribution.
@@
Thermometry, the process of estimating the temperature of a sample, plays a crucial role in quantum thermodynamics Campbell et al. (2017); De Pasquale and Stace (2018); Zhang and Tong (2022) and has various applications, such as determining macroscopic objects in their ground state Chan et al. (2011); Montenegro et al. (2020). In the context of previously discussed equilibrium quantum sensing, it has been demonstrated that finite temperature can surprisingly enhance the performance of critical quantum metrology protocols Ostermann and Gietka (2024). In addition, in certain scenarios, local estimation for thermometry may not be sufficient, making an alternative estimation strategy highly important. Note that the bound in Eq. (87) may not be saturated as, in general, the optimal measurement setup vary for different values of the unknown parameter $\lambda$. In some cases that the optimal measurement is not a function of $\lambda$ the bound can indeed be saturated. An example of such scenario is thermometry for which energy measurement is optimal, independent of the temperature of the system. This allows to design an optimal probe for measuring temperature over an arbitrary interval. In Ref. Correa et al. (2015), it is shown that optimization of the probe reduces to optimizing the energy eigenvalues. In Ref. Campbell et al. (2018), this analysis was extended to investigate the spectral features required for sensitivity to multiple temperatures. Interestingly, for the case of local thermometry, where temperature varies over a narrow interval, the optimal probe contains only two distinct eigenenergies, with a unique ground state and full degeneracy for the rest of the spectrum. The energy gap between the ground state and the rest of the spectrum depends on the temperature Paris (2015)and can be obtained by solving a nonlinear equation Correa et al. (2015). Using the global sensing formulation of Eq. (87), one can show that the optimal probe requires more distinct energy eigenvalues as the temperature varies over a larger interval Mok et al. (2021). In fact, through exploiting an evolutionary algorithm the authors of Ref. Mok et al. (2021) find out that an effective two-level probe for local sensing turns into a three-level system with the two lowest energy eigenvalues being unique and then the rest of the spectrum gets fully degenerate as the temperature interval increases. By further enhancing the interval more and more energy levels are separated from the degenerate band. It is worth emphasizing that the bounds, given by Eq. (87) can be saturated for all these optimal thermometers.
@@
Figure 20: Global sensing for transverse Ising quantum many-body probe $(a)$ Fitting coefficients $a$ and $b$ as a function of $\Delta\lambda_{z}/J$. A smooth transition from the Heisenberg limit ($b=2$) to the standard limit ($b=1$) is observed. Notably, for a prior distribution as wide as $0\leq\Delta\lambda_{z}/J\lesssim 0.07$, the global sensor still delivers quantum-enhanced sensitivity ($b>1$). $(b)$ Minimizing the average uncertainty $g(B_{z})$ always results in better sensitivity when compared to a probe without applying control fields. Figure taken from Montenegro et al. (2021).
@@
Figure 21: Global sensing for temperature estimation: $(a)$ Log-log plot of the global optimum $\bar{\epsilon}_{\mathrm{opt}}$ and local Cramér–Rao-like bound (dot-dashed blue) for a gas of $n$ non-interacting spin-1/2 particles in thermal equilibrium. The global optimum is lower than the local bound unless $n\rightarrow\infty$, indicating that the local bound misses information for small $n$. $(b)$ Data analysis in global thermometry. The global estimate converges to the true temperature after $\mu\simeq 10^{2}$ shots ($\mu$ are energy measurements on the $n$-spin gas). In contrast, the local theory leads to a biased estimate even for $\mu\simeq 500$. Figure taken from Rubio et al. (2021).
@@ XV.1 Optimal Control Based Approaches
Quantum optimal control theory (QOCT) is now an area of extensive literature. We refer the interested reader to the review in Ref. Koch and Budich (2022) for a more detailed treatment of QOCT in other areas of quantum technology. For sensing purposes within QOCT approach, an extra control Hamiltonian is added to the probe’s evolution. If the probe evolves through a Hamiltonian $H_{\lambda}$ containing information about the unknown parameter $\lambda$, then the time-dependent total Hamiltonian is given by
$H_{tot}(t)=H_{\lambda}+H_{c}(\textbf{u}(t),t),$ (93)
where the control Hamiltonian $H_{c}$ is a function of control parameters $\textbf{u}(t)$. The goal is to design the control Hamiltonian $H_{c}(\textbf{u}(t),t)$ in such a way that maximum precision in estimation of the parameter $\lambda$ is achieved. This can be broadly accomplished in two ways: (quasi)analytically and numerically. In terms of control strategies, broadly two distinct strategies are possible, viz., open-loop control and closed-loop, i.e., feedback control.
@@ XV.1.1 Open-loop QOCT strategies
The simplest and oldest systematic approach to QOCT is to apply Pontryagin’s Maximum Principle (PMP), which builds on the calculus of variations to derive a set of necessary conditions for optimal control, see Ref. Boscain et al. (2021) for a detailed introduction to PMP in quantum settings. PMP introduces an auxiliary system described in terms of a costate $|\mu(t)\rangle$ which must satisfy the original evolution equation backwards in time, i.e., $i\hbar\frac{|d\mu(t)\rangle}{dt}=-H_{tot}(t)|\mu(t)\rangle$ with the same boundary conditions. Now if the task is to optimize a cost-function $\mathcal{C}$, which can be generally written as a sum of two cost functions, i.e., $\mathcal{C}=\mathcal{C}_{1}+\mathcal{C}_{2}$, with $\mathcal{C}_{1}$ being a cost function in terms of some suitable path-dependent quantity integrated over the trajectory of evolution (say, the average QFI for sensing tasks) and $\mathcal{C}_{2}$ being a terminal cost function (say, the fidelity achieved with a target state etc); then PMP puts the necessary condition for the control parameters being optimized, i.e., $\textbf{u}(t)=\textbf{u}^{\text{opt}}(t)$, as the optimization over the following function for every time $t$ and the quantum state $|\psi(t)\rangle$ along the original evolution equation
$\textbf{u}^{\text{opt}}(t)=\arg\max_{\textbf{u}(t)}\left[\langle\mu(t)|H_{tot}|\psi(t)\rangle-\mathcal{C}\right].$ (94)
For very simple problems, this optimization is sometimes possible to perform analytically. In Ref. Lin et al. (2021), the authors consider such a problem, where the twist and turn angular momentum Hamiltonian $H=\chi J_{z}^{2}+\omega J_{z}$ with unknown parameter $\omega$ is augmented with a control Hamiltonian $\Omega(t)J_{x}$ and find the optimal trajectory of $\Omega(t)$ by maximizing the QFI of the final state as the cost function. However, for more complicated problems like those involving many-body interactions, analytically solving Eq. (94) is no longer presumed to be viable. Thus numerical algorithms must be adopted. Popular gradient-based optimization algorithms for quantum control include gradient ascent pulse engineering (GRAPE) Khaneja et al. (2005) and Krotov’s method Reich et al. (2012). In the former, updates are made simultaneously for all times $t$, while in the latter updates are made sequentially. However, for exponentially growing Hilbert space dimensions in generic many-body systems, gradient-based algorithms are highly limited. Moreover, in noisy environments gradient updates are highly unreliable in the first place. Thus gradient-free optimization algorithms are often required. Primary approach along this direction is to Fourier-decompose the control pulse sequence into a smaller number of bases. Randomly sampling a small number of bases via chopped random basis (CRAB) Caneva et al. (2011) thus avoids the dimensionality problem for many-body systems.
@@ XV.1.2 Closed-loop QOCT strategies
The second approach is the so called feedback-based strategies, where measurement results obtained in each round are fed into the quantum control for the next round. Pang and Jordan Pang and Jordan (2017) considered a phase estimation problem for a parameter $\lambda$ imprinted into a time-dependent Hamiltonian $H(\lambda,t)$. The goal is to optimize the QFI along the entire trajectory. To do so, they noted that QFI at each instant $T$ is upper bounded by the following integral
@@
where $\nu_{\max}(t)$ and $\nu_{min}(t)$ are respectively maximum and minimum eigenvalues of $\partial_{\lambda}H(\lambda,t)$ at any time $t$. Thus, Pang and Jordan employed an additional control Hamiltonian $H_{c}(t)$ to steer the trajectory towards the eigenspace of $\partial_{\lambda}H(\lambda,t)$ as much as possible. The crucial aspect is that the control Hamiltonian is independent of the actual parameter $\lambda$, but written in terms of $\tilde{\lambda}$, a known prior estimate of $\lambda$ which is updated at each round. As an illustration, they considered a qubit in a uniformly rotating magnetic field $\textbf{B}(t)=B(\cos\omega t~{}\hat{x}+\sin\omega t~{}\hat{y})$, where the goal is to estimate $B$ or the field rotation frequency $\omega$. For estimating $\omega$ from evolution up to time $T$, their approach furnishes a QFI $I_{Q}(\omega)=B^{2}T^{4}$. In contrast, in the absence of the control Hamiltonian, QFI scales only quadratically with time. This result indicates that adaptive control strategies are eminently useful for quantum sensing.
Yang et al in Ref. Yang et al. (2020) addressed a slightly different question, preparation of the optimal probe for maximum precision via adaptive feedback control. They used purity loss instead of QFI for ease of direct measurement. Their approach circumvents the problem of exponentially growing Hilbert spaces by avoiding any computation involving matrix representation of the time-independent Hamiltonian $H$. In particular use for many-body sensing, they considered a spin-chain probe with the control being limited to a sequence of local rotations and even the rotation strength can be adjusted only a fixed number of times. Their results, experimentally implemented on an NMR platform, indicate that Heisenberg scaling is possible even with such a severe limitation on the control.

### v3

@@ XIV Global Estimation Theory: Parameter Estimation with Minimal Prior Information
In contrast to local estimation theory, global estimation theory applies when no (or some) prior information is available. Recently, the impact of uncertainty in control parameters on the sensitivity of critical sensors has been explored Mihailescu et al. (2025a). Therefore, in these situations the unknown parameter varies within a broader sensing interval. Consequently, the choice of measurement to be performed on the probe and the analysis of the estimation data should be optimal on average, that is it should be equally effective for any possible value of the parameter Helstrom (1976); Holevo (2011a).
One of the earliest studies on global estimation theory, aimed at overcoming the limitations of local estimation theory, was proposed by van Trees Trees (1968); Gill and Levit (1995) (a stronger type of inequality is also provided by Klaassen Klaassen (1989)). The van Trees inequality addresses the comprehensive concept of global estimation, where the unknown parameter $\lambda$ varies according to some prior distribution $z(\lambda)$, determining that the average variance is given by the Bayesian Cramér-Rao bound or posterior Cramér-Rao bound Trees (1968); Paris (2009):
@@
Note that, in general, van Trees bound is not tight and cannot be saturated. To achieve a tighter bound a systematic and general metric evaluating the performance of global estimation and its significance for quantum many-body probes with criticality has been examined in Ref. Montenegro et al. (2021). The global estimation procedure involves considering a modified quantum Cramér-Rao bound, which introduces the average uncertainty of the estimation as follows:
@@
where $z(\lambda)$ is the prior distribution of the unknown parameter $\lambda$ to be sensed, $M$ is the number of measurements, $I_{Q}(\lambda|\bm{B})$ is the QFI, $\bm{B}{=}(B_{1},B_{2},\ldots)$ are external tunable parameters interacting with the probe, and the unknown parameter varies over a sensing interval
@@
$\hat{H}=J\sum_{i=1}^{L}\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+1}^{x}-\sum_{i=1}^{L}(B_{z}+\lambda_{z})\hat{\sigma}_{i}^{z},$ (93)
@@
where $L$ is the system size, $\hat{\sigma}_{i}^{x,y,z}$ represents the Pauli operator at site $i$, $J>0$ denotes the exchange interaction, $B_{z}$ corresponds to the controllable magnetic field that can be adjusted, $\lambda_{z}$ stands for the field to be estimated, and periodic boundary conditions are imposed. Assuming that the unknown parameter varies over a wide region $\Delta\lambda_{z}$, the protocol introduced in Ref. Montenegro et al. (2021) systematically optimizes the quantum many-body probe to deliver the best sensitivity performance on average. In the case of single-parameter estimation, the control field $B_{z}$ acts as an offset of the critical point, see Fig. 20(a). However, when the phase diagram becomes more complex or in the presence of multi-parameter estimation, such an offset becomes non-trivial, see Fig. 20(b).
@@
Figure 20: Sketch of phase diagram and the role of global estimation: $(a)$ The global estimation scheme tunes the quantum probe to operate optimally on average. This procedure ensures that the control fields will optimize the probe to exploit the critical point. $(b)$ As the phase diagram becomes more complex, the probe’s offset becomes non-trivial. In general, the minimization of $g(\bm{B})$ (see Eq. (90)) systematically optimizes the probe to exploit criticality across the entire phase diagram. Figure taken from Montenegro et al. (2021).
@@
By defining the size of the probe $L$ as the relevant sensing resource, it has been demonstrated that the QFI scales super-linearly with $L$ over a moderate wide interval of $\Delta\lambda_{z}$ Montenegro et al. (2021). Fig. 21(a) shows the fitting coefficients $a$ and $b$ obtained from the fitting function $g(B_{z}^{*},\Delta\lambda_{z})\sim aL^{-b}+c$, as a function of the interval $\Delta\lambda_{z}/J$. Here, $\Delta\lambda_{z}/J{\rightarrow}0$ represents the case of local estimation sensing, where much prior information about the unknown parameter is at hand. As the figure illustrates, the $b$ coefficient decreases from the (expected) Heisenberg limit ($b=2$) to the standard quantum limit ($b=1$) as the interval $\Delta\lambda_{z}$ increases. The global estimation protocol still demonstrates quantum-enhanced sensitivity for sensing intervals $0\leq\Delta\lambda_{z}\lesssim 0.07J$. In Fig. 21(b), even under the seemingly standard quantum limit (i.e., $b=1$), the global estimation strategy still delivers superior sensing precision compared to a non-optimized probe. This superiority arises from the global estimation strategy, which ensures probe optimization by minimizing $g(\bm{B})$ over the prior distribution.
@@
Thermometry, the process of estimating the temperature of a sample, plays a crucial role in quantum thermodynamics Campbell et al. (2017); Pasquale and Stace (2018); Zhang and Tong (2022) and has various applications, such as determining macroscopic objects in their ground state Chan et al. (2011); Montenegro et al. (2020). In the context of previously discussed equilibrium quantum sensing, it has been demonstrated that finite temperature can surprisingly enhance the performance of critical quantum metrology protocols Ostermann and Gietka (2024). In addition, in certain scenarios, local estimation for thermometry may not be sufficient, making an alternative estimation strategy highly important. Note that the bound in Eq. (90) may not be saturated as, in general, the optimal measurement setup vary for different values of the unknown parameter $\lambda$. In some cases that the optimal measurement is not a function of $\lambda$ the bound can indeed be saturated. An example of such scenario is thermometry for which energy measurement is optimal, independent of the temperature of the system. This allows to design an optimal probe for measuring temperature over an arbitrary interval. In Ref. Correa et al. (2015), it is shown that optimization of the probe reduces to optimizing the energy eigenvalues. In Ref. Campbell et al. (2018), this analysis was extended to investigate the spectral features required for sensitivity to multiple temperatures. Interestingly, for the case of local thermometry, where temperature varies over a narrow interval, the optimal probe contains only two distinct eigenenergies, with a unique ground state and full degeneracy for the rest of the spectrum. The energy gap between the ground state and the rest of the spectrum depends on the temperature Paris (2015)and can be obtained by solving a nonlinear equation Correa et al. (2015). Using the global estimation formulation of Eq. (90), one can show that the optimal probe requires more distinct energy eigenvalues as the temperature varies over a larger interval Mok et al. (2021). In fact, through exploiting an evolutionary algorithm the authors of Ref. Mok et al. (2021) find out that an effective two-level probe for local sensing turns into a three-level system with the two lowest energy eigenvalues being unique and then the rest of the spectrum gets fully degenerate as the temperature interval increases. By further enhancing the interval more and more energy levels are separated from the degenerate band. It is worth emphasizing that the bounds, given by Eq. (90) can be saturated for all these optimal thermometers.
@@
Figure 21: Global estimation protocol for transverse Ising quantum many-body probe $(a)$ Fitting coefficients $a$ and $b$ as a function of $\Delta\lambda_{z}/J$. A smooth transition from the Heisenberg limit ($b=2$) to the standard quantum limit ($b=1$) is observed. Notably, for a prior distribution as wide as $0\leq\Delta\lambda_{z}/J\lesssim 0.07$, the global sensor still delivers quantum-enhanced sensitivity ($b>1$). $(b)$ Minimizing the average uncertainty $g(B_{z})$ always results in better sensitivity when compared to a probe without applying control fields. Figure taken from Montenegro et al. (2021).
@@
Figure 22: Global estimation procedure for temperature estimation: $(a)$ Log-log plot of the global optimum $\bar{\epsilon}_{\mathrm{opt}}$ and local Cramér–Rao-like bound (dot-dashed blue) for a gas of $n$ non-interacting spin-1/2 particles in thermal equilibrium. The global optimum is lower than the local bound unless $n\rightarrow\infty$, indicating that the local bound misses information for small $n$. $(b)$ Data analysis in global thermometry. The global estimate converges to the true temperature after $\mu\simeq 10^{2}$ shots ($\mu$ are energy measurements on the $n$-spin gas). In contrast, the local theory leads to a biased estimate even for $\mu\simeq 500$. Figure taken from Rubio et al. (2021).
@@ XV.1 Optimal Control Based Approaches
Quantum optimal control theory (QOCT) is now an area of extensive literature. We refer the interested reader to the review in Ref. Koch and Budich (2022) for a more detailed treatment of QOCT in other areas of quantum technology. For sensing purposes within QOCT approach, an extra control Hamiltonian is added to the probe’s evolution. If the probe evolves through a Hamiltonian $\hat{H}_{\theta}$ containing information about the unknown parameter $\theta$, then the time-dependent total Hamiltonian is given by
$\hat{H}_{tot}(t)=\hat{H}_{\theta}+\hat{H}_{c}(\textbf{u}(t),t),$ (96)
where the control Hamiltonian $\hat{H}_{c}$ is a function of control parameters $\textbf{u}(t)$. The goal is to design the control Hamiltonian $\hat{H}_{c}(\textbf{u}(t),t)$ in such a way that maximum precision in estimation of the parameter $\theta$ is achieved. This can be broadly accomplished in two ways: (quasi)analytically and numerically. In terms of control strategies, broadly two distinct strategies are possible, viz., open-loop control and closed-loop, i.e., feedback control.
@@ XV.1.1 Open-loop QOCT strategies
The simplest and oldest systematic approach to QOCT is to apply Pontryagin’s Maximum Principle (PMP), which builds on the calculus of variations to derive a set of necessary conditions for optimal control, see Ref. Boscain et al. (2021) for a detailed introduction to PMP in quantum settings. PMP introduces an auxiliary system described in terms of a costate $|\mu(t)\rangle$ which must satisfy the original evolution equation backwards in time, i.e., $i\hbar\frac{|d\mu(t)\rangle}{dt}=-\hat{H}_{tot}(t)|\mu(t)\rangle$ with the same boundary conditions. Now if the task is to optimize a cost-function $\mathcal{C}$, which can be generally written as a sum of two cost functions, i.e., $\mathcal{C}=\mathcal{C}_{1}+\mathcal{C}_{2}$, with $\mathcal{C}_{1}$ being a cost function in terms of some suitable path-dependent quantity integrated over the trajectory of evolution (say, the average QFI for sensing tasks) and $\mathcal{C}_{2}$ being a terminal cost function (say, the fidelity achieved with a target state etc); then PMP puts the necessary condition for the control parameters being optimized, i.e., $\textbf{u}(t)=\textbf{u}^{\text{opt}}(t)$, as the optimization over the following function for every time $t$ and the quantum state $|\psi(t)\rangle$ along the original evolution equation
$\textbf{u}^{\text{opt}}(t)=\arg\max_{\textbf{u}(t)}\left[\langle\mu(t)|\hat{H}_{tot}|\psi(t)\rangle-\mathcal{C}\right].$ (97)
For very simple problems, this optimization is sometimes possible to perform analytically. In Ref. Lin et al. (2021), the authors consider such a problem with the twist and turn angular momentum Hamiltonian $\hat{H}=\chi\hat{J}_{z}^{2}+\omega\hat{J}_{z}+\Omega(t)\hat{J}_{x}$, where the unknown parameter $\omega$ is augmented with a control Hamiltonian $\Omega(t)\hat{J}_{x}$, which is also the turn term. Ref. Lin et al. (2021) then finds the optimal trajectory of $\Omega(t)$ by maximizing the QFI of the final state with respect to $\omega$ as the cost function. However, for more complicated problems like those involving many-body interactions, analytically solving Eq. (97) is no longer presumed to be viable. Thus numerical algorithms must be adopted. Popular gradient-based optimization algorithms for quantum control include gradient ascent pulse engineering (GRAPE) Khaneja et al. (2005) and Krotov’s method Reich et al. (2012). In the former, updates are made simultaneously for all times $t$, while in the latter updates are made sequentially. However, for exponentially growing Hilbert space dimensions in generic many-body systems, gradient-based algorithms are highly limited. Moreover, in noisy environments gradient updates are highly unreliable in the first place. Thus gradient-free optimization algorithms are often required. Primary approach along this direction is to Fourier-decompose the control pulse sequence into a smaller number of bases. Randomly sampling a small number of bases via chopped random basis (CRAB) Caneva et al. (2011) thus avoids the dimensionality problem for many-body systems.
@@ XV.1.2 Closed-loop QOCT strategies
The second approach is the so called feedback-based strategies, where measurement results obtained in each round are fed into the quantum control for the next round. Pang and Jordan Pang and Jordan (2017) considered a phase estimation problem for a parameter $\theta$ imprinted into a time-dependent Hamiltonian $\hat{H}(\theta,t)$. The goal is to optimize the QFI along the entire trajectory. To do so, they noted that QFI at each instant $t$ is upper bounded by the following integral
@@
where $\nu_{\max}(\tau)$ and $\nu_{\min}(\tau)$ are respectively maximum and minimum eigenvalues of $\partial_{\theta}\hat{H}(\theta,\tau)$ at any time $\tau$. Thus, Pang and Jordan employed an additional control Hamiltonian $\hat{H}_{c}(t)$ to steer the trajectory towards the eigenspace of $\partial_{\theta}\hat{H}(\theta,t)$ as much as possible. The crucial aspect is that the control Hamiltonian is independent of the actual parameter $\theta$, but written in terms of $\tilde{\theta}$, a known prior estimate of $\theta$ which is updated at each round. As an illustration, they considered a qubit in a uniformly rotating magnetic field $\textbf{B}(t)=B(\cos\omega t,\sin\omega t,0)$, where the goal is to estimate $B$ or the field rotation frequency $\omega$. For estimating $\omega$ from evolution up to time $t$, their approach furnishes a QFI $I_{Q}(\omega)=B^{2}t^{4}$. In contrast, in the absence of the control Hamiltonian, QFI scales only quadratically with time. This result indicates that adaptive control strategies are eminently useful for quantum sensing. Yang et al in Ref. Yang et al. (2020) addressed a slightly different question, preparation of the optimal probe for maximum precision via adaptive feedback control. They used purity loss instead of QFI for ease of direct measurement. Their approach circumvents the problem of exponentially growing Hilbert spaces by avoiding any computation involving matrix representation of the time-independent Hamiltonian $\hat{H}$. In particular use for many-body sensing, they considered a spin-chain probe with the control being limited to a sequence of local rotations and even the rotation strength can be adjusted only a fixed number of times. Their results, experimentally implemented on an NMR platform, indicate that Heisenberg scaling is possible even with such a severe limitation on the control.
Until now, we assumed that the control itself is precisely known beforehand. This assumption can be relaxed by modelling the control with an extra unknown nuisance parameter, say $\theta^{\prime}$, in contrast to the parameter $\theta$ of interest. Using Eq. (18) with the weight matrix $\mathcal{W}{=}\text{diag}(1,0)$, we can immediately derive that the variance of our estimation for the parameter of interest, namely $\theta$, should satisfy
$\delta\theta^{2}\geq\frac{1}{I_{\theta\theta}\left(1-\frac{I_{\theta\theta^{\prime}}I_{\theta^{\prime}\theta}}{I_{\theta\theta}I_{\theta^{\prime}\theta^{\prime}}}\right)}.$ (99)
This clearly shows that the imperfection of control induces a separate error in estimation. The specifics of this correction will depend on the details of the imperfect control, e.g., whether it originates from calibration error or noise. It will be an interesting question to consider on a case-by-case basis whether this noise outweighs any potential advantage of the many-body approach discussed in this review vis-a-vis interferometric approaches.